# Uncertainty-Guided Optimization on Large Language Model Search Trees

### v2

@@ 3.4 Decision making
$\hat{x}_{j}=\begin{cases}x_{j}&\text{if }x_{j}\text{ is a leaf,}\\ \underset{x_{c}\in\mathrm{children}(x_{j})}{\operatorname*{arg\,max}}a_{j}(x_{c})&\text{otherwise}.\end{cases}$
@@
Note that they have the same costs since the posterior samples for both strategies are readily available due to the backup step below, i.e., we do not actually perform the recursion (2) in this step. Furthermore, the acquisition function (1) can be extended further, e.g. to take into account a penalty term (Section 5.4). In any case, this selection process is done until an unexpanded node is selected.
2. Expansion Given an unexpanded node $x_{i}$, we query the LLM to obtain the top-$k$ most likely children and their corresponding likelihoods. These likelihoods are new observations and we combine them with the prior samples to obtain $n$ posterior samples $(v_{x_{c}})_{n}$ of each child $x_{c}$. This step is the most expensive part of any tree-search algorithm since querying the LLM is costly. Akin to BO, ULTS strives to reduce the number of node expansions and thus minimize the number of LLM calls.
@@
3. Backup We recursively propagate the newly obtained posterior samples $\{(v_{x_{c}})_{n}\}_{c}$ back up the tree until the root via the path selected in the previous steps. This is done to update the posterior samples contained in each node of the path. They will then influence the selection process in the next iteration, updating the exploration-exploitation tradeoff. Different update strategies can be used depending on the choice of the acquisition function (1). When posterior samples of the children node $x_{c}$ are used to compute $a_{i}(x_{c})$, then we propagate up the posterior samples of the newly expanded nodes as in when computing the prior, i.e. by recursively taking their maximum. When the posterior samples of the best descendant of $x_{c}$ are used in $a_{i}(x_{c})$, we simply propagate up the posterior samples of the best child among the newly expanded nodes, without taking further maximums along the path.
@@
Termination criterion  Finally, the posterior samples over the optimal values can not only be used for the selection of new nodes but also to monitor the progress of the optimization. For instance, one can compute the following empirical probability ^$\hat{\mathbb{P}}(c^{*}<v_{\boldsymbol{x}_{0}})=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\left[c^{*}\leq(v_{\boldsymbol{x}_{0}})_{n}\right]$, which corresponds to the probability of the current best likelihood $c^{*}$ among all leaves ULTS has visited so far is lower than the best value $v_{\boldsymbol{x}_{0}}$ at the root, according to our posterior belief. Note that the computation of such a probability is done as in the acquisition function $a(x)$, i.e., using the posterior samples of the root node $\boldsymbol{x}_{0}$ or the posterior samples of the best descendant. Then, one can decide to stop the search once this probability is below some confidence level $\varepsilon>0$.
@@ 3.5 Remarks
Practical considerations  In order to put a tractable upper bound on the runtime of ULTS, we introduce a hyperparameter $k_{\max}$ on the maximum number of nodes that can be expanded per level, similar to beam search or top-$k$ sampling.
Moreover, we stop the search as soon as a set of $k_{\max}$ leaves is attained or the termination probability exceeds below $1-\varepsilon$. We shall see in Section 5 that even under these further constraints, we still obtain good results while being efficient.
Limitations   There are strong reasons to assume that the iid. assumption might not fully hold.
Indeed, our modeling assumptions are chosen with practicality in mind. This is akin to how priors in Bayesian neural networks (Wilson & Izmailov, 2020) are chosen, which are usually simply just isotropic Gaussians. Even in standard BO, a generic Gaussian-process prior/kernel is often assumed by default (Balandat et al., 2020). For a further discussion of the prior assumptions, see Appendix A. Moreover, our work does not take the miscalibration of LLM softmax outputs into account. Indeed, it has been shown that higher likelihood is not always correlated with human preferences (Holtzman et al., 2019; Wiher et al., 2022). However, LLM calibration is orthogonal to the focus of the present paper—we leave this for future work. Nevertheless, ULTS is extensible: One can construct a different prior or use a different utility function to address this; see Section 5.4 for example.
@@ 4 Related Work
Probabilistic/Bayesian optimization on trees has been proposed for game trees. Hennig et al. (2010) developed a roll-out based probabilistic tree search algorithm for game trees, e.g. for playing Tic-Tac-Toe and Go. Crucially, the structure of the game tree is different than the LLM search tree. Moreover, they assumed a Gaussian process prior with the Brownian motion kernel in conjunction with the expectation propagation algorithm (Minka, 2001) to model their beliefs. Grosse et al. (2021) extended their approach for a more general directed acyclic graph structure. In contrast, we focus on the tree implied by the sequential generation process in LLMs with a Dirichlet or empirical prior.
Our method can be seen as utilizing a best-first search strategy, of which the A* algorithm (Hart et al., 1968) is the most famous. Moreover, an A*-like beam search algorithm, under the name of best-first beam search, has also been proposed for decoding LLMs (Meister et al., 2020). Unlike the non-probabilistic A and best-first beam search, we approach optimization-on-tree problems via the lens of decision-making under uncertainty—putting a prior belief about the unknown, updating it based on the observations, and making decision based on the posterior belief. Furthermore, best-first beam search (Meister et al., 2020) has different goals to ours: it is designed to output the same set of leaves (and thus the optimal likelihood value) as the standard beam search. ULTS, meanwhile, focuses on both performance (i.e., attaining higher likelihood than beam search) and efficiency (reducing the number of costly LLM forward passes).
Finally, Monte Carlo tree search, which also utilizes samples to make decisions, has also been proposed for decoding LLMs (Leblond et al., 2021; Liu et al., 2023; Hao et al., 2023; Feng et al., 2023; Zhang et al., 2023; Zhou et al., 2024; DeLorenzo et al., 2024). Different from our goal, they focused on incorporating external rewards that are only observable at the leaf nodes. For instance, Zhang et al. (2023) defined the reward to be the unit-test results for code-generating LLMs. Moreover, Monte-Carlo tree search does not define a probabilistic model (i.e. a prior and posterior beliefs) over quantities in the search tree—its decision making is based on the statistics obtained by exploring the search tree on the fly. In contrast, ULTS does not require costly gathering of those statistics during the decoding process; instead, it makes decisions based on precomputed samples from its beliefs.
@@ 5 Experiments
Setup  In this section, we evaluate ULTS in both close-ended and open-ended text generation problems. We set ULTS’ $k_{\max}\in\{2,3,4,5,10,20\}$ and set $\epsilon$ to a default value of $0.1$. The error bars indicate ±$\pm 1$ standard error of the mean based on all test sentences. We compare ULTS against beam search (the de-facto decoding method for machine translation), as well as the recent baselines multinomial beam search (Multinomial BS; Kool et al., 2019), contrastive search Su et al. (2022) nucleus sampling (Holtzman et al., 2019), best-of-k sampling (Stiennon et al., 2020), speculative decoding (Leviathan et al., 2023), and DoLA (Chuang et al., 2024). For beam search, multinomial beam search, and best-of-k baselines, we evaluate beam sizes and numbers of samples $k\in\{1,2,3,4,5,10,20\}$, respectively. We set any other hyperparameters of all baselines as suggested by the original papers or by Huggingface. We use Huggingface’s implementation of the baselines and provide our implementation of ULTS in the supplements.
Choice of the prior  We sampled a set $\mathcal{C}$ of Categorical distributions from the LLM on training sequences for each of the datasets and both of the LLMs. We used $1000$ samples from the training set for each of the datasets and LLMs we considered. They are used for the empirical prior, as well as training data for fitting the concentration parameter $\alpha$ of the Dirichlet prior.
Figure 7 in Section C.3 shows samples for the maximum of Categorical distributions returned by the LLM, as well as the distribution of the maximum of the Categorical distribution from a Dirichlet prior for $\alpha=10^{-1},10^{-4},5\times 10^{-6}$. For the experiments below, we picked a value for $\alpha$ based on the histograms in Figure 7. For a comparison between different values of the concentration parameter, please refer to Fig. 8 (Section C.3). Note that this way of choosing the hyperparameter $\alpha$ is very convenient compared to performing costly cross-validation.
5.1 Close-ended generation
We compare ULTS with the baselines in a machine translation task with 1000 randomly sampled sequences from the WMT-19 English to German dataset in Fig. 3 in terms of likelihood and BLEU score. For the LLM, we use T5-large (Raffel et al., 2020); for ULTS, we use a Dirichlet prior with $\alpha=5\times 10^{6}$. Unless specified otherwise, we use the strategy in (2) for the selection and backup steps. See Appendix C for results with the other strategy and further dataset/setting. All experiments are done on single NVIDIA GeForce RTX 2080 Ti and NVIDIA A40 48GB GPUs for GPT-2 and Llama-2-7b, respectively. We stop the exploration of a path in the tree the <EOS> token is found or when the maximum depth of 60 is exceeded.
The results are in Figs. 3 and 3. ULTS are both more efficient (fewer node expansion) and more performant (higher average log-likelihood) at all values of $k$/$k_{\text{max}}$. For all methods, small to medium beam sizes of $k=3$ seem to work best in terms of BLEU scores—these scores decrease for a larger (maximum) beam with, in contrast to the log-likelihood in Fig. 3 . See Section C.2, for qualitative examples of the translated sentences.
5.2 Open-ended generation

### v3

@@ 2 Setting
$\Delta_{i}=\begin{cases}1&\text{if $s_{i}$ is a leaf,}\\ \underset{s_{j}\in\mathrm{children}{s_{i}}}{\max}\{\boldsymbol{c}_{i}(a_{j})\cdot\Delta_{j}\}&\text{otherwise.}\end{cases}$ (1)
Example: Decoding Large Language Models  Here, the set of actions $\mathcal{A}$} is a vocabulary consisting of (natural) language tokens. The root of the search tree is given by the context, e.g., a question in a question-answering system. The reward function $R$ is determined by a forward pass through a large language model (LLM), i.e. a large-scale neural network that utilizes the attention mechanism (vaswani2017attention, ) to predict the probability of the next token appearing in a sequence. This process is done recursively until termination; either when a specified depth $D$ has been reached or when a special token $\langle\text{EOS}\rangle$” is selected. Considering the vocabulary size $B$ ranges from around 3232k to 256256k (radford2019language, ; chowdhery2023palm, ) and the reward evaluation amounts to a costly forward pass, even when the sequence length $D$ is small, computing the optimal sequence using standard tree-search algorithms like A or Monte Carlo tree is search infeasible. Thus, in practice, one is limited to using heuristic algorithms that are close to greedy, e.g. beam search and its variants (vijayakumar2016diverse, ; meister2021determinantal, ; freitag2017beam, ).
Beyond LLM decoding, several MDPs fit within our problem formulation, where stepwise likelihood serves as a reward. In general, ULTS is applicable to any autoregressive model, not just in the language domain, but for example also in chemistry tasks (christofidellis2023unifying, ; gao2024generative, ). Further specific example applications include optimizing process parameters for machines (such as paper drying machines in (chen2025reinforcement, )), feature selection (luo2019autocross, ), various neural combinatorial optimization problems by decoding pointer networks (vinyals2015pointer, ), structure learning tasks (kok2005learning, ), and finding Maximum A Posteriori sequences as in (pellom2002efficient, ).
3 Method
Here, we introduce Uncertainty-guided Likelihood-Tree Search (ULTS) which places a prior belief over the log-likelihood reward function and computes the implied posterior samples over nodes’ values. We introduce and discuss the modeling assumptions in section˜3.1.1 and show how to derive approximate beliefs over the optimal values in the search tree in section˜3.1.2. The samples of the posterior beliefs are then used to make decisions about which subtree to expand next and when to stop the search (section˜3.2).
3.1 Probabilistic model
3.1.1 Prior beliefs over immediate rewards
A straightforward belief one can consider is the Dirichlet distribution. For tractability, we assume that the probabilities are iid. draws from a symmetric Dirichlet distribution with parameter $\alpha>0$, i.e., $p(\boldsymbol{c}_{s})=\mathrm{Dir}(\alpha)$. Thus, $\alpha$ controls how concentrated the sampled probability vectors are. For small $\alpha$, one would typically strongly favor a few children, whereas for large $\alpha$ the discrete distribution would closely resemble a uniform distribution over the children. The symmetry of the prior implies in our context that we do not have a preference for particular children a priori, but it is also possible to replace the symmetric prior with an asymmetric one with no additional costs.
In settings where the Categorical distributions are predicted by a trained sequence model $f$, an alternative option is to define an empirical prior over $\boldsymbol{c}_{i}$ based on predictions $\boldsymbol{c}_{i}=f(s_{i})$ on a subset of the training/validation data. Let $\smash{\{s_{0,n}\}_{n=1}^{N}}$ be $N$ samples of root nodes from this subset. We can obtain a set $\{s_{0,n},s_{1,n},\dots,s_{D,n}\}_{n=1}^{N}$ of $D$-step completions of $s_{0,n}$’s with $f$ (e.g. through a greedy search). We can then collect samples of the Categorical distributions from this generation process: $\mathcal{C}=\{\boldsymbol{c}_{0,n},\dots,\boldsymbol{c}_{D-1,n}\}_{n=1}^{N}$. Then, instead of sampling from $\mathrm{Dir}(\alpha)$, we can sample from $p(\boldsymbol{c}_{i})=\mathrm{Unif}(\mathcal{C})$. This prior is more flexible than the Dirichlet prior since no symmetry nor unimodality assumptions are made at the cost of more overhead. In any case, all these priors are precomputed, so they incur a fixed $\mathcal{O}(1)$ cost.
3.1.2 Prior beliefs over values
Having picked a prior for the categorical distributions, we can derive the implied priors over the optimal values $v_{s_{i}}$ for each node $s_{i}$ in the tree (hennig2010coherent, ). As introduced in Section 2, we write the optimal value of a node as $v_{s_{i}}=\log c_{s_{0}\rightarrow s_{i}}+\log\Delta_{i}$. Due to the iid. assumption above (Section 3.1.1), we have the joint distribution $p(c_{s_{0}\rightarrow s_{i}},\Delta_{i})=p(c_{s_{0}\rightarrow s_{i}})\,p(\Delta_{i})$. While all these quantities are not analytically available, we shall show that sampling from the posterior belief $p(v_{s_{i}}\mid c_{s_{0}\to s_{i}})$ is easy if we are able to sample from $p(\Delta_{i})$. Let us, therefore, derive an approximate sampling scheme for $p(\Delta_{i})$. It follows hennig2010coherent ; grosse2021probabilistic who used Gaussian priors. We recursively approximate the prior distribution of the $\Delta_{i}$’s at level $l$ with Beta distributions $\text{Beta}_{l}(\Delta_{i})$. In a bottom-up approach, we generate $N$ samples $\{\max_{j}\boldsymbol{c}_{n}(a_{j})\mid\boldsymbol{c}_{n}\sim p(\boldsymbol{c})\}_{n=1}^{N}$ for a $\Delta_{i}$ starting at level l=D1l=D-1. Using these samples, we empirically fit the parameters of the Beta distribution $\mathcal{B}_{D-1}(\Delta_{i})$ via maximum likelihood. The distributions of the $\Delta_{i}$ are the same for all nodes on the same level due to the iid. assumption, so this has to be done only once. Note that we need this approximation since the distribution of the maximum $\max_{j}\boldsymbol{c}_{i}(a_{j})$ has no known analytic solution. We then continue by recursively sampling sets of the form (one per level) $\{\max_{j}\boldsymbol{c}_{n}(a_{j})\cdot\Delta_{j}\mid\boldsymbol{c}_{n}\sim p(\boldsymbol{c}),\Delta_{j}\sim\mathcal{B}_{l+1}(\Delta)\}_{n=1}^{N}$ for a $\Delta_{i}$ of the level $l$ and using it to fit the parameters of $\mathcal{B}_{l}(\Delta_{i})$. The time complexity is $\mathcal{O}(DBN)$ for computing the approximations, i.e., it is linear in the depth and width of the tree. We emphasize that they can be pre-computed before the search and reused across different decoding runs. Alg. 1 shows the pseudocode.
1:Input: Depth of the tree $D$, branching size of tree $B$, prior $p(\boldsymbol{c})$ over categorical rewards, number of samples used for the approximation NN
2:Output: Table of parameters $\text{params}_{l}$ for Beta distributions $\mathcal{B}_{l}$ for each level $l$ of the tree
3:for $l\leftarrow D-1$ to $1$ do
4:  for $n\leftarrow 1$ to $N$ do
5:   Sample $\boldsymbol{c}_{n}\sim p(\boldsymbol{c})$
6:   for $j\leftarrow 1$ to $B$ do
7:    if l=D1l=D-1 then
8:     $\Delta_{nj}\leftarrow 1$
9:    else
10:     Sample $\Delta_{nj}\sim\mathcal{B}_{l+1}(\Delta\mid\text{params}_{l+1})$
11:    end if
12:   end for
13:   $\Delta_{n}\leftarrow\max\limits_{j=1,\dots,B}(\boldsymbol{c}_{nj}\cdot\Delta_{nj})$
14:  end for
15:  $\text{params}_{l}\leftarrow\texttt{beta-MLE}(\{\Delta_{n}\}_{n=1}^{N})$
16:end for
Algorithm 1 Prior belief over $\Delta$
3.1.3 Posterior beliefs on frontier nodes
Notice that whenever a new node $s_{i}$ is added to the search tree, the likelihood associated with the path from the root to that node $c_{s_{0}\rightarrow s_{i}}$ is fully observed—its distribution is simply a Dirac delta. This means, the joint distribution becomes $p(\boldsymbol{c}_{s_{0}\to s_{i}},\Delta_{i})=\delta(c_{s_{0}\to s_{i}})\,p(\Delta_{i})$. Therefore, to sample $v_{s_{i}}$ given that we have observed $c_{s_{0}\to s_{i}}$—i.e., sampling from the posterior $p(v_{s_{i}}\mid c_{s_{0}\rightarrow s_{i}})$—it is sufficient to sample from $p(\Delta_{i})$ and then simply scale all samples by $c_{s_{0}\rightarrow s_{i}}$. During the search, these samples are backed up the tree and leveraged to make decisions in exploring the tree. Notice that these samples are cheap to get, unlike empirical samples based on simulations/rollouts.
3.2 Algorithm
At each iteration, we use the posterior samples by following the steps of Monte Carlo tree search, but without the expensive rollout step. Pseudocode in appendix˜D.
Step 1: Selection  Starting from the root node $s_{0}$, we recursively pick a child based on an acquisition function until an unexpanded node is selected. From a decision-theoretic perspective, this is done by choosing a utility function that encodes our preferences about the outcome (e.g. high likelihood) and then integrating out the unknown variables influencing this outcome. Let $\boldsymbol{v}_{i}$ contain the (unknown) optimal values of the descendants of $s_{i}$ and let $u(s_{c},\boldsymbol{v}_{i})$ be a utility function. The idea is to select the child node $s_{c}$ that maximizes the expected utility:
$\int u(s_{c},\boldsymbol{v}_{i})\,p(\boldsymbol{v}_{i}\mid c_{v_{s_{0}\to s_{i}}})\,d\boldsymbol{v}_{i},$ (2)
For ULTS, we use an utility function that encodes our preference for finding a descendant with high optimal value: $u(s_{c},\boldsymbol{v}_{i}):=\mathbb{I}[v_{\hat{s}_{c}}=\max_{s_{j}\in\mathrm{children}(s_{i})}v_{\hat{s}_{j}}]$. Different realizations of $\hat{s}_{c}$ and $\hat{s}_{j}$ can be used: The most straightforward is to use the children’s optimal values themselves. This corresponds to simply setting $\hat{s}_{j}=s_{j}$ and $\hat{s}_{c}=s_{c}$.
Another strategy is to use the optimal values of each child’s best descendant—intuitively, it ignores other paths in the tree that might have higher values, leading to a more pessimistic belief. This can be beneficial to bias the search towards more exploration (for further discussion, see section˜C.7). In this case, $\hat{s}_{j}$ is defined as (similarly for $\hat{s}_{c}$):
$\hat{s}_{j}=\begin{cases}s_{j}&\text{if }s_{j}\text{ is a leaf,}\\ \underset{s_{c}\in\mathrm{children}(s_{j})}{\operatorname*{arg\,max}}a_{j}(s_{c})&\text{otherwise}.\end{cases}$
@@ 3.2 Algorithm
Note that they have the same costs since beliefs over the optimal values for both strategies are readily available due to the backup step below, i.e., we do not actually perform the recursion (3) in this step.
The acquisition function is then derived as a sample-based approximation to the expectation in (2):
[…]
Note that we can replace the above utility function with any other utility function and then use the posterior samples to derive the acquisition function (wilson2018maximizing, ). For example, we can use a utility function that encodes a preference for sequences with fewer repetition, see section˜C.9.
Figure 1: An example of two iterations with ULTS. The upper row show the observed categorical distirbution (left), the implied posterior over the optimal values in log space (center) and the resulting acquisition function over the children in the first level of the tree (right). The lower two rows show the corresponding quantities for the first and second level of the tree after the second iteration.
Step 2: Expansion  Given an unexpanded node $s_{i}$, we evaluate its children’s rewards. These newly obtained rewards are new observations that can be combined with the prior samples to obtain $N$ posterior samples $\{v_{s_{c},n}\}_{n=1}^{N}$ of each child $s_{c}$. In turn, they will affect the acquisition function in the next iteration.
This step is often the most expensive part of any tree-search algorithm since we assume costly evaluation of the rewards. Fewer node expansions is thus preferable.
@@
Step 3: Backup  We recursively propagate the newly obtained posterior samples $\{v_{s_{c},n}\}_{n=1}^{N}$ back up the tree until the root via the path selected in the previous steps. This is done to update the posterior samples contained in each node of the path. They will then influence the selection process in the next iteration, updating the exploration-exploitation tradeoff.
Different update strategies can be used depending on the choice of the acquisition function (4). When posterior samples of the children node $s_{c}$ are used to compute $a_{i}(s_{c})$, then we propagate up the posterior samples of the newly expanded nodes as in when computing the prior by recursively taking their maximum, i.e. the $n$-th sample for the optimal value of parent node $s_{p}$ is given by the maximum over the $n$-th sample of the children’s optimal values:
$v_{s_{p},n}=\max_{s_{c}\in\mathrm{children}{(}s_{p})}v_{s_{c},n}.$ (5)
When the posterior samples of the best descendant of $s_{c}$ are used in $a_{i}(s_{c})$, we simply propagate up the posterior samples of the best child among the newly expanded nodes, without taking further maximums along the path:
$v_{s_{p},n}=v_{s_{c}*,n},\quad\text{where }s_{c}^{*}=\underset{s_{c}\in\mathrm{children}(s_{p})}{\operatorname*{arg\,max}}a_{p}(s_{c}).$ (6)
Figure˜1 shows examples for the first two iterations of ULTS under the former acquisition function and backup scheme. See section˜C.7 for the corresponding illustration for the latter.
@@
Step 4: Termination  The posterior samples over the optimal values can not only be used for the selection of new nodes but also to monitor the progress of the optimization. For instance, one can compute the following empirical probability $\hat{\mathbb{P}}(c^{*}<v_{s_{0}})=\frac{1}{N}\sum_{n=1}^{N}\mathbb{I}\left[c^{*}\leq v_{s_{0},n}\right]$, which corresponds to the probability of the current best likelihood $c^{*}$ among all leaves ULTS has visited so far is lower than the best value $v_{s_{0}}$ at the root, according to our posterior belief.
Note that the computation of such a probability is done as in the acquisition function $a(x)$, i.e., using the posterior samples of the root node $s_{0}$ or the posterior samples of the best descendant. Then, one can decide to stop the search once this probability is below some confidence level $\varepsilon>0$.
@@ 3.3 Remarks
Practical considerations  In order to put a tractable upper bound on the runtime of ULTS, we introduce a hyperparameter $k_{\max}$ on the maximum number of nodes that can be expanded per level, similar to beam search. Moreover, we stop the search as soon as a set of $k_{\max}$ leaves is attained or the termination probability exceeds below $1-\varepsilon$. We shall see in section˜5 that even under these further constraints, we still obtain good results while being efficient.
Limitations  In this work we focus on the uncertainty arising from the exponentially growing search space under limited compute resources. However, there might be other sources of uncertainty, e.g. uncertainty about the likelihood/reward function $R$. In LLM decoding, there is evidence that higher likelihood is not always correlated with human preferences (holtzman2019curious, ; wiher2022decoding, ; stahlberg2019nmt, ; eikema2020map, ; zhang2020trading, ). Appendix C contains additional evaluations of ULTS in terms of other metrics than the likelihood, where there is indeed no perfect correlation. There is an ongoing line of research on how to counteract miscalibration in LLMs that is complementary to our work yang2023bayesian ; kapoor2024large ; onal2024gaussian ; tonolini2024bayesian , but combining and evaluating such approaches with ULTS goes beyond the scope of this paper. Moreover, the iid. prior assumption might not fully hold, but in our inference scheme it is essential to obtain tractability. This is akin to how priors in Bayesian neural networks (wilson2020bayesian, ) are usually chosen, which are just iid. Gaussians.
@@ 4 Related Work
Probabilistic tree search methods have been explored for various applications. hennig2010coherent ; tesauro2012bayesian introduced a method for game trees like Go, using a Gaussian process prior and expectation propagation (minka2001family, ) to model the beliefs. grosse2021probabilistic extended this to directed acyclic graphs. mern2021bayesian also used Gaussian processes for general planning problems. Unlike these methods, we focus on trees with a Categorical likelihood reward function. Additionally, the above methods rely on MCTS-based roll-outs, which do not scale efficiently. ULTS avoids costly rollout statistics by leveraging precomputed belief samples for decision-making.
Our approach employs a best-first search strategy, akin to A* (hart1968astar, ). An A*-like beam search algorithm, under the name of best-first beam search, has also been studied in meister2020best . Unlike the non-probabilistic A and best-first beam search, we approach optimization-on-tree problems via the lens of decision-making under uncertainty—putting a prior belief about the unknown, updating it based on the observations, and making decision based on the posterior belief.
MCTS is a staple uncertainty-aware search method and it has also been proposed for various LLM applications (leblond2021machine, ; liu2023don, ; hao2023reasoning, ; feng2023alphazero, ; zhang2023planning, ; zhou2023language, ; delorenzo2024make, ). However, unlike ULTS, MCTS is non-probabilistic and generally costly since it requires rollouts. Moreover, it is often used for cases where rewards are only observable at leaf nodes (zhang2023planning, , etc.) or when external value estimators is available (feng2023alphazero, , etc.). The latter introduces a substantial overhead in evaluating a value network, which is often another large-scale model. ULTS, meanwhile, uses a simple model which introduces negligible overhead (see fig.˜8).
Minimum Bayes risk methods (kumar2004minimum, ; eikema2021sampling, ; bertsch2023itsmbrwaydown, ) also rely on the maximization of an expected utility function. However, they do not employ probabilistic models and do not aim to make decisions in a sequential manner. Rather, they generate a set of full-sequence samples and make a decision on which sequence in the set to be selected.
@@ 5 Experiments
figures
In this section, we evaluate ULTS in both on-model and off-model problems. In the former, we assume that the rewards are sampled from a Dirichlet distribution. In the latter, we use black-box LLMs as reward functions. Unless specified otherwise, we use the acquisition function in (3) for the selection and backup steps. See appendix˜C for results with the other strategy and further dataset/setting. All experiments are done on a single NVIDIA GeForce RTX 2080 Ti and NVIDIA A40 48GB GPUs for GPT-2 and Llama-2-7b, respectively. Errorbars reported in the plots and tables indicate the standard error of the mean as calculated with scipy.stats.sem.
5.1 On-model experiments
We compare ULTS to beam search on artificially generated search problems from Dirichlet priors. The trees have branching factor $B=8$ and depth $D=5$. Since these trees are so small we optimized the acquisition function in (4) over the entire frontier, i.e. non-recursively with $k_{\text{max}}=\infty$ like standard A*. The transition probabilities at each node are sampled from a Dirichlet prior with fixed $\alpha\in\{0.1,0.2,0.5,0.8\}$. The comparison is on-model, i.e., ULTS is run with the ground truth parameter of $\alpha$. We repeat the experiment with different values for the confidence parameter $\varepsilon$ of ULTS from $\{0.05,0.1,0.3,0.5\}$. Beam search is run with beam sizes ranging from $1$ to $7$. Since the $\Delta$-Terms of ULTS play the same role as the heuristic terms in A*, we also add A* with uninformative, optimistic heuristic (i.e. estimating the probability of the remaining sequence parts with 1) for an additional layer of ablation. The results in fig.˜2 and fig.˜5 in Appendix C.1 show that ULTS dominates across the entire range of hyper-parameters. This suggests that knowledge about concentration strength helps reduce the number of search steps.
Figure 2: Comparison of ULTS with beam search and A* on synthetic rewards sampled from a Dirichlet distribution with $\alpha=0.2$ and $\alpha=0.8$. ULTS finds the optimum efficiently. Errorbars indicate one times the standard error of the mean.
5.2 Off-model experiments with LLMs